OpenStudy (anonymous):

Can someone jus show me how to solve? NO DRAMA PLEASE Solve: −3x − 3 = 12 5 −5 −3 3

OpenStudy (anonymous):

-3 times a -5 would give you positive 15. Positive 15 minus 3 is 12. Your answer would be the second one. (-5)

OpenStudy (anonymous):

-3x -3=12 -3x=12+3 -3x=15 -x=15/3 -x=5 x=-5

OpenStudy (anonymous):

Thanks :)